OpenStudy (anonymous):

find derivative of (secx)/(1+cosx) please show steps.

OpenStudy (anonymous):

\[\frac{(1+cosx)\times (secx)' - secx ( 1 +cosx)'}{(1 + cosx)^2}\]

OpenStudy (anonymous):

(secx)' = secx * tanx (1 + cosx)' = -sinx

OpenStudy (anonymous):

yeah im still stuck after that. I tried factoring out sec

OpenStudy (anonymous):

secx*tanx(1 +cosx) + secx*sinx --------------------------- (1 + cosx)^2 secx(tanx + tanx*cosx + tanx) -------------------------- (1 + cosx)^2 secx(2tanx + sinx) ------------------ (1 + cosx)^2 2tanx*secx + tanx --------------- (1 +cosx)^2 I guess this much simplification must be enough

OpenStudy (anonymous):

uhm on the second step what happened to sinx?

OpenStudy (anonymous):

sinx*secx = sinx*1/cosx = sinx/cosx =tanx

myininaya (myininaya):

ishaan didn't you factor the sec(x) out though?

OpenStudy (anonymous):

yeah thats what i was thinking to myinanaya

myininaya (myininaya):

\[(\frac{\sec(x)}{1+\cos(x)})'=\frac{(\sec(x))'(1+\cos(x))-\sec(x)(1+\cos(x))')}{(1+\cos(x))^2}\] \[=\frac{\sec(x)\tan(x)(1+\cos(x))-\sec(x)(0-\sin(x))}{(1+\cos(x))^2}\] \[=\frac{\sec(x)\tan(x)+\sec(x)\tan(x)\cos(x)+\sec(x)\sin(x)}{(1+\cos(x))^2}\] \[=\frac{\sec(x)\tan(x)+\tan(x)+\tan(x)}{(1+\cos(x))^2}\] \[=\frac{\sec(x)\tan(x)+2\tan(x)}{(1+\cos(x))^2}\] \[=\frac{\tan(x)(\sec(x)+2)}{(1+\cos(x))^2}\]

OpenStudy (anonymous):

on your thrid step, what happened to secxtanxcosx? cause a tan is replaced there

myininaya (myininaya):

\[=\frac{\frac{\sin(x)}{\cos(x)}(\frac{1}{\cos(x)}+2)}{(1+\cos(x))^2}\] \[=\frac{\frac{\sin(x)}{\cos^2(x)}+2 \cdot \frac{\sin(x)}{\cos(x)}}{(1+\cos(x))^2} \cdot \frac{\cos^2(x)}{\cos^2(x)}\] \[=\frac{\sin(x)+2 \sin(x) \cos(x)}{\cos^2(x)(1+\cos(x))^2}=\frac{\sin(x)(1+2\cos(x))}{1+\cos(x))^2}\]

myininaya (myininaya):

\[\sec(x)=\frac{1}{\cos(x)}\] right?

myininaya (myininaya):

so \[\frac{1}{\cos(x)} \cdot \cos(x)=?\]

OpenStudy (anonymous):

1

myininaya (myininaya):

so what is 1*tan(x)?

OpenStudy (anonymous):

tanx -___- okay I see. THANKS U SO MUCH!. do u think u can help me on one more please?

myininaya (myininaya):

i have to go to bed its late and i have to wake up at 5 am :(

OpenStudy (anonymous):

awesome latex

myininaya (myininaya):

ishaan can help right?

OpenStudy (anonymous):

I will try goodnight myininaya

myininaya (myininaya):

you will do fine but you did make one error above did you see it?

OpenStudy (anonymous):

good night myininaya!

myininaya (myininaya):

goodnight and good luck you guys! I believe in you!

OpenStudy (anonymous):

error hmm no I didn't what is it?

myininaya (myininaya):

well i see one error and i stop reading after that

myininaya (myininaya):

ok i'm going for real

OpenStudy (anonymous):

okay so what is the derivative of (8x-4)(2x^-2+x^-4)

OpenStudy (anonymous):

(8x-4)'*(2x^(-2) + x^(-4)) + (2x^(-2) + x^(-4))'*8x-4 you can do it

OpenStudy (anonymous):

I actually tried for 2x^-2+x^-4 I got for the derivative -4/x^3 -4/x^5...is that correct?

OpenStudy (anonymous):

oh yeah I see it now @myininaya thanks

OpenStudy (anonymous):

it is correct, good job

OpenStudy (anonymous):

okay so I have (8x-4)((5x^7+1)+(2x^-2+x^-4)(8) correct?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

then i have: (-32/x^3+16x^5)+(16x^-2+8x^-4)

OpenStudy (anonymous):

except (8x -4)(5x^7 +1) part how'd you get this

OpenStudy (anonymous):

oh my bad it was suppose to be times -4/x^3 -4/x^5

OpenStudy (anonymous):

you are right then

OpenStudy (anonymous):

(-32/x^3+16x^5)+(16x^-2+8x^-4) correct too right?

OpenStudy (anonymous):

(8x -4)(-4/x^3 -4/x^5) = -32/x^2 - 32/x^4 + 16/x^3 + 16/x^5 -32/x^2 - 32/x^4 + 16/x^3 + 16/x^5 + (16/x^2+8/x^4)

OpenStudy (anonymous):

reply to '(-32/x^3+16x^5)+(16x^-2+8x^-4) correct too right?' check it again doesn't seem correct to me

OpenStudy (anonymous):

yeah i forgot to foil and put the x's on the denominator. okay so after that, what happens to the x denominators?

OpenStudy (anonymous):

-32/x^2 - 32/x^4 + 16/x^3 + 16/x^5 + (16/x^2+8/x^4) -16/x^2 - 24/x^4 + 16/x^3 + 16/x^5

OpenStudy (anonymous):

uhm how is it -24/x^4?

OpenStudy (anonymous):

oh wiat nvm

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

no problem